### v1

# Contextual Embedding-based Clustering to Identify Topics for Healthcare Service Improvement

Abstract
Understanding patient feedback is crucial for improving healthcare services, yet analyzing unlabeled short-text feedback presents significant challenges due to limited data and domain-specific nuances. Traditional supervised learning approaches require extensive labeled datasets, making unsupervised methods more viable for uncovering meaningful insights from patient feedback. This study explores unsupervised methods to extract meaningful topics from 439 survey responses collected from a healthcare system in Wisconsin, USA. A keyword-based filtering approach was applied to isolate complaint-related feedback using a domain-specific lexicon. To delve deeper and analyze dominant topics in feedback, we explored traditional topic modeling methods, including Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM), alongside BERTopic, an advanced neural embedding-based clustering approach. To improve coherence and interpretability where data are scarce and consist of short-texts, we propose kBERT, an integration of BERT embeddings with k-means clustering. Model performance was assessed using coherence scores ($C_{v}$) for topic interpretability and average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$) for topic diversity. Results indicate that kBERT achieves the highest coherence ($C_{v}$ = 0.53) and distinct topic separation ($\operatorname{IRBO_{avg}}$ = 1.00), outperforming all other models in short-text healthcare feedback analysis. Our findings emphasize the importance of embedding-based techniques for topic identification and highlight the need for context-aware models in healthcare analytics.
Index Terms:
LDA, GSDMM, BERT, BERTopic, k-means, Clustering, Topic Modeling, Complaint Identification, Healthcare, Short-text
publicationid: pubid: copyright-placeholder
I Introduction
In recent years, online forums have become pivotal in recognizing complaints about services or products. Linguistic studies categorize a complaint as a primary form of communication for expressing dissatisfaction when expectations are not met concerning a product, situation, organization, or event[1]. Both complaints and compliments hold significant weight in business decisions, with complaints being particularly crucial. The analysis of customer product reviews, a practice long utilized in e-commerce, has greatly enhanced business operations and customer satisfaction, ensuring customers feel valued and heard. Similarly, various sectors including banking systems and city utility services are now integrating customer feedback mechanisms to refine their services.
Complaint identification is a nuanced subset of sentiment analysis. Many existing methods have employed supervised machine learning with trained datasets, which are often costly and time-intensive to prepare for domain-specific applications[2]. This challenge has naturally led to an increased interest in unsupervised learning approaches, particularly topic modeling and clustering techniques. These techniques can uncover underlying patterns in textual data without the need for labeled datasets.
Topic modeling aims to identify latent themes within a collection of documents by analyzing word distributions. Methods such as Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM) have been widely applied to textual data to uncover dominant topics [3]. Clustering methods, in contrast, group similar documents based on their feature representations. Unlike topic models, which allow words to belong to multiple topics, clustering techniques assign each document to a single group based on textual similarity. Recent advances in neural embedding-based clustering, such as BERTopic, have been explored for their potential to handle short and noisy texts compared to LDA and GSDMM. [4, 5]. While topic modeling methods and neural embedding-based clustering approaches have been explored for various text analysis tasks, their effectiveness in scenarios where data is both short and limited remains an open question. In this study, we develop a neural embedding-based clustering model, kBERT, and systematically assess its performance alongside established methods to evaluate their suitability for analyzing short-text on limited data.
This study introduces a system designed to identify topics within patient feedback based on surveys from hospital visits. The system aims to highlight complaint comments by analyzing survey feedback, thereby assisting hospital authorities in enhancing service quality and understanding patient needs more effectively.
Our methodology begins with a manually curated list of keywords to segregate positive and negative comments. The keyword list incorporates domain-specific terms that help capture the nuanced details of patient feedback (see Table II). To complement this keyword-based approach and gain deeper insights into the underlying themes in patient feedback, we employ unsupervised learning methods for topic extraction. We evaluate the performance of traditional topic modeling techniques (LDA, GSDMM) and modern embedding-based clustering approaches (BERTopic) in extracting meaningful topics from a small short-text dataset to better understand the applicability of these methods in patient feedback analysis. Additionally, we compare their results with kBERT to assess its potential in addressing the challenges posed by limited and short healthcare-related textual data. These methods allow for a more comprehensive analysis of patient feedback, addressing both latent topic discovery and document similarity-based clustering to improve topic coherence and interpretability.
Given the challenges associated with short-text length, small dataset sizes, and domain-specific issues in patient feedback, this research aims to explore the following research questions (RQs):
RQ1: How can unsupervised techniques be effectively utilized to identify and categorize topics in short-text patient feedback?
RQ2: What are the key challenges in analyzing healthcare-specific short-texts and how can advanced natural language processing (NLP) techniques improve extracting meaningful insights from patient feedback in a hospital setting?
RQ3: How can the performance of unsupervised topic modeling and clustering methods be effectively evaluated in short-text patient feedback analysis?
The rest of the paper is structured as follows: Section II reviews related work on complaint identification, topic modeling, and clustering, highlighting the challenges of short-text analysis in healthcare. Section III details the proposed methodology, including data preprocessing, keyword filtering, and applying unsupervised techniques, addressing RQ1. Section IV presents the result analysis, comparing the performance of both topic modeling and clustering approaches, evaluating their coherence, diversity, and interpretability, thereby addressing RQ3. Section V discusses the findings and their implications for healthcare service improvement, effectively answering RQ2. Finally, Section VI concludes the study.
II Related Work
II-A Topic Modeling and Word Embeddings-based Clustering
In the field of NLP, complaint identification, particularly when labeled datasets are available, is often approached as a classification task. This area employs various machine learning models like support vector classifier (SVC), Random Forest, Decision Tree, and multilayer perception (MLP), along with deep learning classifiers such as Dense Neural Networks and convolutional neural networks (CNNs) [6]. Additionally, NLP libraries like VADER, TextBlob, and NLTK are commonly used to derive sentiment scores, enabling the classification of feedback into positive, negative, or neutral categories [7]. However, these libraries might not always capture nuanced details, particularly in domain-specific contexts. With the rise of Large Language Models (LLMs), the field of complaint classification is evolving. L. Yuan et al. introduced a Complaint Text Classification Framework (CCF) based on LLMs, which leverages advanced feature extraction techniques from large language models, significantly enhancing classification accuracy [8].
While classification alone is useful, it doesn’t provide deep insights into the nature of the text. LDA can be effective in understanding the underlying concepts of language. LDA identifies the latent topics within a text corpus. As demonstrated in subsequent studies, LDA is frequently used for sentiment analysis in complaint data. For instance, A. Karami et al. utilized LDA to categorize GEICO customer complaints into four distinct groups [9], while N. Hu et al. applied LDA and Structural Topic Modeling to analyze hotel reviews in New York City, identifying ten major themes of customer dissatisfaction and how these vary across different hotel grades [10]. M. N. Aziz et al. adopted a mixed approach for opinion classification, leveraging both Lexicon-Based (SentiWordNet) and Machine Learning-Based (TF-IDF with SVM) methods, and employed LDA for topic clustering [11]. As research in this field progresses, various topic modeling methods have been compared to evaluate their effectiveness. For instance, H. S. Jung et al. evaluated four topic modeling methods—LDA, Nonnegative Matrix Factorization (NMF), Combined Topic Models (CTM), and BERTopic. Their findings showed that BERT based model demonstrated superior performance across both topic diversity and coherence [12].
Embeddingbased methods have significantly advanced topic modeling by utilizing contextual embeddings like BERT and SBERT to capture semantic and syntactic nuances. The Embedded Topic Model (ETM) [13] integrates LDA with embeddings to enhance topic coherence. Approaches combining BERT embeddings with clustering methods [14, 15] show improved coherence and interpretability. For instance, a framework integrating BERT with clustering and dimensionality reduction [14] demonstrated superior clustering quality, while a hybrid approach combining BERT, LDA [15] achieved high topic coherence. C-Top2Vec [16] further refined topic modeling through multi-vector representations and hierarchical clustering for granular insights.
Topic modeling has evolved with the introduction of advanced models, yet traditional methods still hold significant advantages in certain applications. GPTopic, a novel approach leveraging LLMs, enhances the interpretability and interaction of topic representations, surpassing the traditional “top-word” approach by allowing dynamic modifications and chat-based interactions with topics. However, GPTopic can occasionally generate inaccurate results due to hallucinations, particularly when applied to datasets with fewer than 10,000 documents [17]. While LLMs, including those used in GPTopic, offer flexibility and zero-shot capabilities, they struggle with maintaining topic granularity and often produce overlapping or redundant topics, which complicates the identification of distinct themes [18]. In contrast, traditional methods like LDA and BERT base model remain consistent and reliable, generating coherent topics without hallucinations, making them highly adaptable across various domains without extensive fine-tuning [19].
However, despite their advancements, LLMs require substantial computational resources, making their deployment costly due to the need for high-performance GPUs or TPUs for fine-tuning and inference. This makes traditional methods like LDA and clustering-based approaches more practical for resource-constrained environments. Additionally, LLM-enhanced applications that provide APIs or interfaces like ChatGPT [20], pose risks of data leakage, as models trained on large datasets can unintentionally expose sensitive information. The lack of clarity in how data is stored, processed, and transmitted further exacerbates trust, security, and privacy risks [21, 22].
TABLE I: Datasets
Dataset Name Size Median Type Availability
Patient Feedback 439 14 Short-text private
Clickbait-title 10,000 9 Short-text public
Clickbait-title (500) 500 9 Short-text public
20Newsgroup 15,021 83 Long-text public
20Newsgroup (500) 500 86 Long-text public
II-B Topic Modeling for Short-Text
Conventional methods like LDA encounter challenges due to their reliance on bag-of-words representations, which can be inadequate for capturing the nuances in short texts[23]. Recent research highlights the effectiveness of alternative models. For instance, GSDMM and BERTopic have shown promising results, outperforming LDA in handling short-text. One study by Udupa et al. [23] specifically explored the use of GSDMM and BERTopic for efficient topic clustering in short-text scenarios. Meanwhile, Jiang et al. [24] implemented the Associated Topic Model (ATM) which leverages a collapsed Gibbs sampling algorithm, further emphasizing the diversity of approaches suitable for short-text analysis. Additionally, clustering token-level contextualized word representations from models like BERT has proven effective in capturing polysemy and organizing documents, performing comparably or better than LDA topic models, as found in recent evaluations [25]. These studies collectively suggest a shift towards models that can better interpret the limited contextual information available in shorter documents.
II-C Topic Modeling in Healthcare Feedback Service
LLMs are increasingly being applied to healthcare-related NLP tasks [26], while recent studies have focused on analyzing patient feedback to enhance healthcare service quality. For instance, Alexander et al. [27] have developed a tool for data analysis and visualization that leverages the BERTopic model to identify key topics and trends in patient feedback collected from NHS services in England. This approach enables a dynamic overview of sentiment changes over time. Similarly, Osváth et al. [28] have applied BERT embeddings coupled with sentiment analysis to dissect patient experiences and public reactions sourced from a Hungarian online forum. Their work aims to uncover meaningful patterns and emotional responses, which are critical for enhancing the quality of healthcare services. These studies illustrate the growing application of advanced topic modeling techniques in healthcare feedback analysis, highlighting their potential to improve service understanding and delivery.
To best of our knowledge, no prior study has addressed short-text topic modeling on limited datasets, specifically in the context of healthcare service.
TABLE II: Sample Keyword (Hot Word) List
infect angry malpractice delay MRSA
blood sepsis cops grievance mis
trauma wound abus argu un
negl viol massac hipaa hippa
african american gun bruise burn
III Methodology
This study focuses on short-text patient feedback, specifically in small datasets, where traditional topic modeling techniques often face challenges. We apply keyword filtering to identify complaint-related feedback using a curated lexicon, ensuring a focused analysis on patient grievances. Simultaneously, topic modeling is employed to extract dominant themes, providing insights into both complaints and general concerns.
As our dataset consists of short-text feedback, GSDMM is used as the baseline for short-text topic modeling, while LDA is included as a reference for long-text topic modeling. These models are compared with BERTopic and kBERT, which leverage transformer-based embeddings to enhance topic coherence in small short-text datasets.
III-A Dataset Description
Patient comments are collected from a health system in Wisconsin, USA, through surveys. This feedback offers detailed insights into patient experiences. The surveys include feedback on various aspects, such as hospital stays, emergency department visits, clinic appointments, ambulatory surgery, outpatient services, and urgent care. A total of 439 patient comments were collected for analysis and testing. To filter complaint comments, the ‘Patient Experience and Quality’ team created a list of 202 specific keywords, referred to as “hot words”, including domain-specific terms related to medical procedures, patient safety, and service quality. The dataset is private and consists of short-text feedback, with a median word count of 14.
To evaluate the robustness of our methodology, we also tested it on two additional public datasets: the Clickbait-title dataset [29] and the 20Newsgroup dataset[30]. The Clickbait-title dataset is a short-text dataset comprising headlines categorized as clickbait or non-clickbait. We selected 10,000 samples from this dataset, with a median word count of 9, making it representative of short-text content. Additionally, to examine model performance on small short-text datasets, we extracted a subset of 500 samples (clickbait-title (500)). In contrast, the 20Newsgroup dataset consists of longer articles from 20 distinct topics. It is a widely used benchmark dataset for different natural language tasks. To standardize the data, we excluded the bottom and top 10% of articles based on word count and retained 15,021 articles for evaluation and the median word count is 83. We also extracted a subset of 500 samples (20Newsgroup (500)) to explore small dataset scenarios in long-text settings.
These datasets allow for a comprehensive evaluation of topic modeling and clustering techniques across varying text lengths (short vs. long) and dataset sizes (large vs. small), providing insights into how different approaches perform under diverse conditions.
Input:
$K=\{k_{1},k_{2},...,k_{202}\}$ // keyword list
$C=\{C_{\text{full}},C_{\text{short}},C_{\text{prefix}},C_{\text{jargon}}\}$ // keyword categories
for $w\in C_{\text{short}}$ do
       if w is ambiguous then
             Refine $w$ to full, meaningful form
       end if
end for
for $k\in C_{\text{short}}$ or $C_{\text{full}}$ do
       // generate all POS forms
       $K_{\text{pos}}\leftarrow Wf(k)=\{N(k),V(k),A(k),R(k)\}$
       $K_{\text{lemma}}\leftarrow$ lemmatize($K_{\text{pos}}$)
end for
Algorithm 1 Keyword Process
III-B Keyword Processing
In this work, we utilize a carefully curated keyword list to effectively filter and categorize the patient feedback into positive and negative comments. These keywords, which are crucial for analyzing the sentiment of the feedback, are presented in Table II. This lexicon ensures a comprehensive reflection of the terminology commonly found in patient complaints, aiding in the accurate assessment of their sentiment. The keyword list is also important as it contains domain-specific nuances. Some feedback includes medical and legal terms that may not explicitly mention complaints. For instance, “I saw another patient’s information on the clinic’s screen, which seems like a HIPAA issue”. These subtle details are often missed by general-purpose models. Capturing these nuances is crucial for improving patient care. In Table II, the keywords are systematically classified into four distinct groups to enhance the precision of complaint detection in the texts. This classification is as follows:
  • Full Keywords: These encompass complete terms that are often used in their entirety within complaints, such as “accident”, “delay”, “agitate”, “american” etc.
  • Shortened Keywords: Truncated forms of words, like “abus” (for abuse) and “argu” (for argue), which are used in different parts of speech (POS) in the feedback.
  • Prefix Keywords: Terms like “mis” and “un”, serve as prefixes, indicative of negative connotations when attached to other words.
  • Jargon/Medical Term Keywords: This category includes essential terms like “sepsis”, “bruise”, “HIPAA”, and “MRSA”, crucial for identifying comments on various topics such as regulatory compliance or infection-related concerns. ‘HIPPA’ is included as it is a frequent misspelling of ‘HIPAA’ by patients.
Each group is designed to address the different ways keywords could manifest in the data — as standalone words, abbreviated forms, or as part of medical jargon — thus providing a comprehensive filter for the preliminary categorization of comments.
In our methodology, the keyword processing phase presented a unique challenge — the ambiguity arising from shortened keywords. For example, the abbreviation ‘compl’ is extracted from terms such as ‘complaint’ and ‘complete’, though only the former is relevant to our analysis. To mitigate this, we refined our keyword list to include only the full, unambiguous forms of each word, ensuring that only contextually relevant terms are considered.
Another complexity arose from the morphological forms of keywords within the dataset. Words like ‘argue’ appeared in various derivatives, such as ‘argues’, ‘argued’, or ‘arguing’. Traditional lemmatization is employed to address this, which involves reducing words to their base or dictionary form. Lemmatization’s effectiveness is contingent upon accurate POS tagging. The process needs to recognize the grammatical role of words within the context of their use. This is because POS tagger models are typically trained on complete sentences or documents rather than isolated tokens. This training approach poses a challenge when a word like ‘argument’ needs to be reduced to ‘argue’ [31].
To navigate these linguistic nuances, we utilize the ‘word_forms’ library in Python [32]. This library provides a comprehensive mapping of word forms across different POS categories: nouns, verbs, adverbs, and adjectives; allowing us to accurately identify and process the various permutations of a keyword in our dataset. The complete keyword processing procedure is detailed in Algorithm 1.
Input:
$F=\{f_{1},f_{2},...,f_{439}\}$ // patient feedback
$S=\{a,\text{and},\text{the},\text{of},\ldots\}$ // stop words
$K_{\text{lemma}}\leftarrow$ from Algorithm 1
for $f\in F$ do
       $T_{f}\leftarrow$ tokenize($f$)
       $R_{f}\leftarrow$ remove $T_{f}$ in ($S$)
       for $t\in T_{f}$ do
             if $t$ starts with $C_{\text{prefix}}$ or $t\in C_{\text{jargon}}$ or $t\in K_{\text{lemma}}$ then
                   $F_{\text{neg}}\leftarrow f$
             end if
            else
                   $F_{\text{pos}}\leftarrow f$
             end if
       end for
end for
Algorithm 2 Feedback Process and Filter
III-C Feedback Preprocessing
The fidelity of topic modeling in NLP is heavily reliant on the quality of input data. Recognizing this, our methodology is built on a thorough protocol for preprocessing data. Initially, tokenization—the segmentation of text into individual terms—is conducted to granulate the dataset into analyzable units.
Subsequent to tokenization, we focus on the elimination of stop words. These words, while structurally integral, contribute minimally to the semantic load of the text. Common stop words such as ‘a’, ‘and’, ‘the’, and ‘of’ are identified and removed. Their high frequency and low informational value make them prime candidates for exclusion to enhance data purity. To perform this task, we employ the ‘spacy’ library—a powerful and versatile Python toolkit for advanced NLP [33]. Its sophisticated algorithms allowed for an efficient and thorough filtration process, excising stop words from the tokenized text. This cleansing step is vital, as it significantly reduces noise within the data, thereby streamlining the path for more accurate and meaningful topic modeling. The detailed procedure for this filtration is presented in Algorithm 2.
III-D Topic Identification
III-D1 Latent Dirichlet Allocation (LDA)
LDA is a generative probabilistic model widely used for uncovering latent topics within a collection of text documents [34]. LDA assumes that each document is a mixture of various topics, where each topic is characterized by a probability distribution over words. By breaking down a document-word matrix into two smaller matrices—the Document-Topic Matrix and the Topic-Word Matrix—LDA reveals the relationships between documents and their underlying topics, as well as the association between topics and the defining words. This process makes LDA particularly valuable for analyzing large collections of text data. The LDA model operates through a hierarchical generative process. At the document level, a topic distribution $\theta$ for each document $d$ is drawn from a Dirichlet distribution with parameter $\alpha$, which serves as the prior for the document-topic distributions. For each word $w_{d,n}$ in a document, a topic $z_{d,n}$ is sampled from the document’s topic distribution $\theta_{d}$. Subsequently, a word is generated by sampling from the topic-specific word distribution $\beta_{z_{d,n}}$, which is drawn from another Dirichlet distribution with parameter $\eta$. This layered structure captures both the document-level topic diversity and the word-level thematic coherence.
Figure 1: LDA Blueprint (Source:[13, 35] )
In our study, we applied LDA to analyze hospital patient feedback documents, identifying and categorizing latent themes within the text. Each document was modeled as a probabilistic mixture of topics, with $\theta_{d,k}$ representing the likelihood of topic $k$ in document $d$, and $\beta_{k,w}$ indicating the relevance of word $w$ within topic $k$. To enhance thematic clarity, we experimented with 3 to 20 topic clusters to balance specificity and breadth. The Dirichlet parameter $\alpha$ was set to ‘auto,’ allowing dynamic adjustment of document-topic distributions. We also fine-tuned the top words per topic, improving the model’s ability to highlight relevant terms and themes.
III-D2 Gibbs Sampling Dirichlet Mixture Model (GSDMM)
Alongside LDA, we also implement the GSDMM for compliant topic identification. The integration of GSDMM is particularly significant given the nature of our data: most survey responses were brief, often comprising 1 to 3 sentences [36]. GSDMM is uniquely suited for such short-texts, as it operates under the assumption that a single document predominantly expresses a single topic. This is a notable deviation from LDA’s approach, where a document can encompass multiple topics, making LDA more suitable for longer texts [37].
The implementation of GSDMM involved modeling each document as a mixture of topics, with a single dominant topic assigned to each document. For a given document $d$, the topic $k$ is assigned based on the conditional probability $P(k\mid d)$, which is calculated using Equation 1. In this equation, $n_{k,-d}$ represents the number of documents assigned to topic $k$, excluding the current document $d$; $N_{-d}$ is the total number of documents excluding $d$; $K$ is the total number of topics; is the count of word $n_{k,w}$ w in topic k; $n_{k}$ is the total number of words assigned to topic $k$; and $V$ is the vocabulary size. The hyperparameters $\alpha$ and $\beta$ serve as priors for the document-topic and topic-word distributions, respectively, analogous to their roles in LDA. Through iterative sampling, GSDMM converged to a stable distribution of topics across the documents, effectively capturing the dominant theme in each text fragment.
$P(k|d)\propto\frac{n_{k,-d}+\alpha}{N_{-d}+K\alpha}\times\frac{n_{k,w}+\beta}{n_{k}+V\beta}$ (1)
III-D3 BERTopic
BERTopic [38] is a topic modeling technique that integrates transformer-based embeddings, dimensionality reduction, clustering, and topic representation to extract meaningful insights from textual data. It begins with a transformer embedding model, such as BERT or RoBERTa, to generate high-dimensional contextual embeddings that capture semantic relationships between words and documents. To enhance computational efficiency and clustering performance, UMAP (Uniform Manifold Approximation and Projection) is applied for dimensionality reduction, preserving essential structures in a lower-dimensional space. The reduced embeddings are then clustered using HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise), which dynamically determines the optimal number of topics based on data density, enabling the discovery of latent topics in an unsupervised manner. This method allows BERTopic to automatically generate topics without requiring the user to specify a predefined number. Additionally, BERTopic framework [39] offers the flexibility to manually specify the desired number of topics, allowing for both unsupervised and user-controlled topic generation. Finally, c-TF-IDF (class-based Term Frequency-Inverse Document Frequency) is employed to extract the most representative words for each cluster, improving topic interpretability. Combining these four components, BERTopic provides a scalable and flexible approach to topic modeling, making it particularly effective for analyzing large, unstructured text data. However, BERTopic has some limitations. It assumes each document belongs to a single topic, which may not fully capture multi-topic documents.
III-D4 k-means clustering with BERT embedding (kBERT)
To address the unique challenges posed by our small dataset of short-texts, we developed a customized topic modeling pipeline that integrates a customized pre-processing approach, BERT embeddings and k-means clustering. Our topic identification system leverages the Bidirectional Encoder Representations from Transformers (BERT), a transformer-based model pre-trained on a vast corpus of text, to generate rich, contextualized embeddings. These embeddings are particularly effective in capturing the subtle nuances of short-text data, making them ideal for our application. Additionally, BERT embeddings can effectively handle Out-of-Vocabulary (OOV) words, making them particularly useful for dealing with domain-specific terminology. Importantly, we only utilized the embedding layer of BERT while freezing all other layers. This ensured that the model acted purely as a feature extractor, preventing any modifications to its pre-trained weights. As a result, there is no risk of overfitting or catastrophic forgetting, making our approach particularly well-suited for small datasets [40]. We avoid dimensionality reduction methods, as they can remove key linguistic features, especially in low-resource settings where every word holds significant meaning [41]. The high-dimensional embeddings generated by BERT are preserved to retain the full contextual representation.
Our approach began with careful preprocessing of the textual data. Stopwords were removed, and the text was lemmatized to ensure uniformity. Additionally, we extended the stopword list with domain-specific and custom terms to improve data quality. Each preprocessed comment was then encoded into embeddings using the ‘bert-base-uncased’ [^1] https://huggingface.co/google-bert/bert-base-uncased variant of the pre-trained BERT model. This model was chosen for its balance between performance and computational efficiency, as well as its ability to generalize well across diverse datasets. The uncased variant ensured the focus remained on content, minimizing case-sensitive variations, which is particularly advantageous for short-texts.
The generated embeddings served as input for a k-means clustering algorithm, where we experimented with different numbers of clusters ($n$ = 3, 5, 10 etc) to find the optimal grouping for the data. For topic interpretability, we extracted the top $k$ words (e.g., 5,10) from each cluster by analyzing word frequencies in representative comments. To improve topic interpretability, we first selected a set of representative comments for each cluster based on their proximity to the cluster centroid in the BERT embedding space. These comments were then analyzed to extract the most frequently occurring words, ensuring that the topic words accurately reflected the cluster’s semantic meaning. This iterative process allowed us to refine the topic selection and ensure the clusters were meaningful and cohesive.
IV Result Analysis
IV-A Feedback Filtering
Our dataset, encompassing 439 comments, underwent keyword matching, which served as a quantitative measure of the filtering process’s effectiveness. Notably, the matching frequency of keywords within the comments indicates a substantial impact on the accuracy of the filtering mechanism. Initial results without the use of lemmatization presented 89 keyword matches. The introduction of lemmatization, a process facilitating the reduction of words to their base or ‘lemma’ form, slightly increased this match count to 92.
A further improvement is achieved by generating and matching keywords across different POS after lemmatization. This approach propelled the match count to a peak of 102 instances. This increment underscores the pivotal role of nuanced linguistic techniques in enhancing the sensitivity of our keyword-based feedback-filtering approach.
[…]
Dataset Model Number of Topics
5 10 20 30 40
Clickbait-title LDA 0.35 0.37 0.35 0.34 0.28
GSDMM 0.44 0.47 0.49 0.48 0.47
BERTopica 0.38 0.46 0.44 0.43 0.46
kBERT 0.40 0.39 0.41 0.49 0.48
20Newsgroup LDA 0.57 0.51 0.55 0.52 0.54
GSDMM 0.46 0.49 0.57 0.53 0.58
BERTopicb 0.41 0.38 0.39 0.41 0.43
kBERT 0.36 0.39 0.40 0.47 0.43
IV-B Topic Modeling
In their work, R. Egger et al. highlight the importance of selecting the appropriate model based on the nature of the data and emphasize the need for careful interpretation of results [42]. Building on this perspective, our topic modeling efforts focused on finding the ideal balance between coherence and clarity. To ensure a comprehensive evaluation, we assessed our models’ performance using two metrics: the widely used coherence score ($C_{v}$) for topic interpretability and Rank-Biased Overlap (RBO) for evaluating topic diversity.
Coherence score $C_{v}$ quantitatively assesses the coherence of topics. A higher $C_{v}$ score indicates better performance, suggesting that the top words of a topic are not only frequent but also semantically similar. This indicates a clearer and more meaningful topic representation. The $C_{v}$ measure is derived from the Normalized Pointwise Mutual Information (NPMI) [43]. It accounts for the probability of words co-occurring within topics relative to their individual probabilities. The NPMI tends to be particularly attuned to less common words and is well-suited for analyzing small datasets like ours [44]. It can be calculated using the equation 2 [45]. In this equation, $\mathrm{P}\left(w_{i}^{r},w_{i}^{s}\right)$ is the joint probability of the words occurring together, while $\mathrm{P}\left(w_{i}^{r}\right)$ and $\mathrm{P}\left(w_{i}^{s}\right)$ are the individual probabilities of the words. The epsilon $\epsilon$ is a small number added to prevent division by zero. The $C_{v}$ is defined by equation 3. $\gamma$ is a parameter that can be tuned to give more importance to higher PMI values. This tuning allows for a weighted contribution of word pairs based on their PMI.
$\operatorname{NPMI}\left(w_{i}^{r},w_{i}^{s}\right)=\frac{\log_{2}\frac{P\left(w_{i}^{r},w_{i}^{s}\right)+\epsilon}{P\left(w_{i}^{r}\right)P\left(w_{i}^{s}\right)}}{-\log_{2}\left(P\left(w_{i}^{r},w_{i}^{s}\right)+\epsilon\right)}$ (2)
$\mathrm{C}_{\mathrm{V}}\left(w_{i}^{r},w_{i}^{s}\right)=\mathrm{NPMI}^{\gamma}\left(w_{i}^{r},w_{i}^{s}\right)$ (3)
To evaluate the effectiveness of different topic modeling approaches, we computed the Coherence Score (Cv) for each model across three datasets: Clickbait-title, 20Newsgroup, and Patient Feedback. The datasets were divided into large and small subsets to assess model robustness across different data scales. Given that the Patient Feedback dataset consists of short-text comments, we tested our models on public short-text datasets for validation. For the analysis, we primarily focus on short-text, especially for small datasets, to evaluate how well our models handle limited data. Coherence scores are measured using the top 10 topic words for large-text datasets and the top 5 topic words for short-text datasets. The best results for short-text are marked boldface in Tables III and IV.
For large datasets, we generated topics at different levels (5, 10, 20, 30, and 40 topics). As shown in Table III, GSDMM consistently outperformed other models in short-text datasets, achieving the highest coherence score of 0.49 for 20 topics in Clickbait-title. This indicates its ability to effectively capture coherent topics in short-texts. Additionally, kBERT also achieved a high coherence score of 0.49 at 30 topics for short-text datasets (Clickbait-title). This motivates further exploration of kBERT for small datasets, where capturing meaningful patterns in limited data is crucial. LDA, while performing well in long-text scenarios, struggled in short-text datasets.
BERTopic, which dynamically determines the number of topics, generated 163 topics in 20Newsgroup (Cv = 0.60) and 178 topics in Clickbait-title (Cv = 0.49). This adaptability allowed BERTopic to excel in large datasets by identifying a diverse range of topics. However, in smaller datasets, it generated a limited number of topics, reducing coherence performance. In the Patient Feedback dataset, BERTopic produced 3 topics with a coherence score of 0.46, indicating moderate effectiveness in short-text scenarios.
For small datasets, we generated 3, 5, 10, 15, and 20 topics. As presented in Table IV, kBERT performed exceptionally well in handling short-text patient feedback, achieving the highest coherence score of 0.53 for 3 topics, demonstrating its strong ability to extract meaningful insights from limited and sparse patient feedback data. Additionally, in the Clickbait-title (500) dataset, kBERT achieved a strong coherence score of 0.45 for 5 topics, reinforcing its strength in short-text small dataset scenarios. This suggests that BERT embeddings are particularly useful for extracting meaningful insights from short and sparse healthcare-related feedback.
[…]
Dataset Model Number of Topics
3 5 10 15 20
Clickbait-title (500) LDA 0.29 0.30 0.31 0.29 0.30
GSDMM 0.28 0.38 0.32 0.37 0.36
BERTopica N/A N/A N/A N/A N/A
kBERT 0.36 0.45 0.45 0.44 0.42
20Newsgroup (500) LDA 0.53 0.49 0.47 0.49 0.44
GSDMM 0.48 0.47 0.43 0.47 0.36
BERTopicb N/A N/A N/A N/A N/A
kBERT 0.41 0.48 0.43 0.43 0.41
Patient Feedback (439) LDA 0.34 0.35 0.34 0.38 0.36
GSDMM 0.49 0.43 0.42 0.46 0.46
BERTopicc 0.46 N/A N/A N/A N/A
kBERT 0.53 0.49 0.48 0.48 0.41
TABLE V: Topics from kBERT with Examples
Topic Name Topic words Example Complain?
Staff Conduct and Cleanliness good, clean, great, everyone, professional Everyone was careful in the emergency department No
Safety Protocols compliance mask, staff, wearing, safe, one The staff’s masks were reassuring in these uncertain times. No
Efficiency of Service doctor, time, appointment, waiting, room Very long wait from the time I entered the room until the doctor arrived Yes
TABLE VI: IRBOavg Score on Top 5 words and 3 Topics
Dataset Model Score
Patient Feedback LDA 0.83
GSDMM 0.97
BERTopic 1.00
kBERT 1.00
Rank-Biased Overlap (RBO) is a similarity measure for ranked lists that emphasizes top-weightedness and accommodates incomplete or non-conjoint lists [46]. It is particularly useful for evaluating topic diversity in topic modeling by comparing the top-ranked terms across topics. The RBO is defined by the equation 4. $A_{d}$ is the agreement at rank $d$, and $p$ ($10<p<$ 1) adjusts the weight of deeper ranks. Higher $p$ values increase the depth considered, ensuring meaningful topic comparisons. RBO scores range from 0 (no overlap) to 1 (identical lists). Adjust the parameter $p$ to focus on higher or deeper ranks. To provide a more comprehensive diversity measure, we used Inverted RBO $(1-\operatorname{RBO})$, where higher values indicate greater topic diversity by reflecting lower overlap in top-ranked terms. We further compute the Averaged Inverted RBO by averaging pairwise inverted RBO values, as defined by equation 5. Here, A, B, and C represent the ranked lists of top words from three different topics, and their pairwise inverted RBO scores quantify the extent of dissimilarity among them. A higher $\operatorname{IRBO_{avg}}$ value indicates greater diversity in the extracted topics.
$\operatorname{RBO}(S,T,p)=(1-p)\sum_{d=1}^{\infty}p^{d-1}\cdot A_{d}$ (4)
$\begin{split}\operatorname{IRBO_{avg}}(A,B,C,p)=\frac{1}{3}\Big{[}(1-\operatorname{RBO}(A,B,p))+\\ (1-\operatorname{RBO}(A,C,p))+\\ (1-\operatorname{RBO}(B,C,p))\Big{]}\end{split}$ (5)
In this study, we selected 3 topics for the Patient Feedback dataset and evaluated topic diversity using Average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$). A higher IRBO score indicates greater topic distinctiveness by measuring lower overlap in top-ranked terms. As shown in Table VI, BERTopic and kBERT achieved the highest diversity score (1.00), suggesting completely distinct topic distributions. GSDMM (0.97) also demonstrated strong topic separation, while LDA (0.83) exhibited comparatively lower diversity, indicating some overlap in identified topics.
V Discussion
Our topic modeling analysis identified three primary areas of concern in patient feedback: ‘Staff Conduct and Cleanliness’, ‘Safety Protocols Compliance’, and ‘Efficiency of Service’. Positive feedback highlighted professionalism and hygiene practices, reinforcing confidence in staff attentiveness. Similarly, the emphasis on safety measures, particularly mask compliance, indicated patient reassurance regarding institutional health protocols. However, the most significant complaint centered on ‘Efficiency of Service’, with frequent mentions of long wait times and appointment delays, emphasizing the need for healthcare facilities to enhance operational efficiency. These key areas of focus are summarized in Table V, which presents the dominant topics along with representative terms and examples.
A crucial step in our approach was keyword filtering, which effectively differentiated complaints from non-complaints within the dataset. By applying a targeted keyword-based filtration method, we identified 102 complaint-related sentences out of 439 total comments.
In evaluating the topic modeling approaches, kBERT consistently demonstrated superior performance in handling short-text complaint identification, achieving the highest coherence score ($C_{v}$ = 0.53 for 3 topics) and an IRBO score of 1.00, indicating distinct and well-separated topics. This suggests that BERT-based embeddings effectively capture the nuanced meaning of patient complaints, making it the most reliable model for healthcare-related short-text analysis. GSDMM also performed well, particularly in short-text datasets, as its cluster-based approach assigns a dominant topic to each document, making it effective for structured short-form feedback. BERTopic, while excelling in larger datasets due to its dynamic topic generation, struggled in smaller datasets, as observed in its moderate coherence score ($C_{v}$ = 0.46) for the Patient Feedback dataset. LDA, a traditional topic modeling approach, showed strong performance in long-text settings but struggled with short-text patient complaints, reaffirming its dependence on a larger corpus for meaningful topic extraction.
These findings emphasize the importance of choosing context-aware models like kBERT for short-text analysis in healthcare settings. The combination of keyword filtering and effective topic modeling ensures that patient concerns are identified and addressed accurately, leading to data-driven improvements in service delivery. The insights gained from this study can help healthcare providers prioritize patient concerns, streamline complaint resolution, and enhance overall service efficiency, ultimately improving the patient experience and quality of care.
VI Conclusion
This study explored unsupervised techniques for complaint topic identification in short-text patient feedback, comparing traditional topic modeling methods (LDA, GSDMM) with advanced neural embedding-based approaches (BERTopic, kBERT). Our results demonstrated that kBERT outperforms other models in both coherence and diversity metrics, making it the most effective method for analyzing short and sparse healthcare-related texts. Additionally, a keyword-based filtering approach significantly improved complaint detection. The identified complaint categories—staff conduct and cleanliness, safety protocol compliance, and efficiency of service—provide actionable insights for healthcare service improvement. In the future, we will also explore few-shot and instruction-based prompting to improve complaint classification and topic extraction in healthcare feedback analysis. To achieve this, we will use open-source LLMs (e.g., LLaMA) locally with sufficient computational resources, preventing data leakage and ensuring compliance.

### v3

@@
 Abstract
-Understanding patient feedback is crucial for improving healthcare services, yet analyzing unlabeled short-text feedback presents significant challenges due to limited data and domain-specific nuances.
-Traditional supervised learning approaches require extensive labeled datasets, making unsupervised methods more viable for uncovering meaningful insights from patient feedback.
-This study explores unsupervised methods to extract meaningful topics from 439 survey responses collected from a healthcare system in Wisconsin, USA.
-A keyword-based filtering approach was applied to isolate complaint-related feedback using a domain-specific lexicon.
-To delve deeper and analyze dominant topics in feedback, we explored traditional topic modeling methods, including Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM), alongside BERTopic, an advanced neural embedding-based clustering approach.
-To improve coherence and interpretability where data are scarce and consist of short-texts, we propose kBERT, an integration of BERT embeddings with k-means clustering.
-Model performance was assessed using coherence scores ($C_{v}$) for topic interpretability and average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$) for topic diversity.
-Results indicate that kBERT achieves the highest coherence ($C_{v}$ = 0.53) and distinct topic separation ($\operatorname{IRBO_{avg}}$ = 1.00), outperforming all other models in short-text healthcare feedback analysis.
-Our findings emphasize the importance of embedding-based techniques for topic identification and highlight the need for context-aware models in healthcare analytics.
-Index Terms:
-LDA, GSDMM, BERT, BERTopic, k-means, Clustering, Topic Modeling, Complaint Identification, Healthcare, Short-text
-publicationid: pubid: copyright-placeholder
+Understanding patient feedback is crucial for improving healthcare services, yet analyzing unlabeled short-text feedback presents challenges due to limited data and domain-specific nuances.
+Traditional supervised approaches require extensive labeled datasets, making unsupervised methods more practical for extracting insights.
+This study applies unsupervised techniques to analyze 439 survey responses from a healthcare system in Wisconsin, USA.
+A keyword-based filter was used to isolate complaint-related feedback using a domain-specific lexicon.
+To identify dominant themes, we evaluated traditional topic models such as Latent Dirichlet Allocation (LDA) and Gibbs Sampling Dirichlet Multinomial Mixture (GSDMM) - alongside BERTopic, a neural embedding-based clustering method.
+To improve coherence and interpretability in sparse, short-text data, we propose kBERT, which integrates BERT embeddings with k-means clustering.
+Model performance was assessed using coherence scores ($C_{v}$) and average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$).
+kBERT achieved the highest coherence ($C_{v}$ = 0.53) and topic separation ($\operatorname{IRBO_{avg}}$ = 1.00), outperforming all other models.
+These findings highlight the value of embedding-based, context-aware models in healthcare analytics.
 I Introduction
-In recent years, online forums have become pivotal in recognizing complaints about services or products.
-Linguistic studies categorize a complaint as a primary form of communication for expressing dissatisfaction when expectations are not met concerning a product, situation, organization, or event[1].
-Both complaints and compliments hold significant weight in business decisions, with complaints being particularly crucial.
-The analysis of customer product reviews, a practice long utilized in e-commerce, has greatly enhanced business operations and customer satisfaction, ensuring customers feel valued and heard.
-Similarly, various sectors including banking systems and city utility services are now integrating customer feedback mechanisms to refine their services.
-Complaint identification is a nuanced subset of sentiment analysis.
-Many existing methods have employed supervised machine learning with trained datasets, which are often costly and time-intensive to prepare for domain-specific applications[2].
-This challenge has naturally led to an increased interest in unsupervised learning approaches, particularly topic modeling and clustering techniques.
-These techniques can uncover underlying patterns in textual data without the need for labeled datasets.
+Understanding and addressing complaints is essential across domains, including healthcare.
+Linguistic studies define a complaint as a form of communication expressing dissatisfaction when expectations are unmet [17].
+While industries like e-commerce and utilities increasingly use feedback to improve services, healthcare systems are also beginning to leverage patient comments for quality enhancement.
+Complaint identification is often approached through supervised learning, but such methods require large labeled datasets, which are resource-intensive to build.
+As a result, unsupervised techniques like topic modeling and clustering have gained attention for their ability to extract meaningful insights from unlabeled text.
 Topic modeling aims to identify latent themes within a collection of documents by analyzing word distributions.
@@
 Recent advances in neural embedding-based clustering, such as BERTopic, have been explored for their potential to handle short and noisy texts compared to LDA and GSDMM.
-[4, 5].
+[8].
 While topic modeling methods and neural embedding-based clustering approaches have been explored for various text analysis tasks, their effectiveness in scenarios where data is both short and limited remains an open question.
 In this study, we develop a neural embedding-based clustering model, kBERT, and systematically assess its performance alongside established methods to evaluate their suitability for analyzing short-text on limited data.
-This study introduces a system designed to identify topics within patient feedback based on surveys from hospital visits.
-The system aims to highlight complaint comments by analyzing survey feedback, thereby assisting hospital authorities in enhancing service quality and understanding patient needs more effectively.
-Our methodology begins with a manually curated list of keywords to segregate positive and negative comments.
-The keyword list incorporates domain-specific terms that help capture the nuanced details of patient feedback (see Table II).
-To complement this keyword-based approach and gain deeper insights into the underlying themes in patient feedback, we employ unsupervised learning methods for topic extraction.
-We evaluate the performance of traditional topic modeling techniques (LDA, GSDMM) and modern embedding-based clustering approaches (BERTopic) in extracting meaningful topics from a small short-text dataset to better understand the applicability of these methods in patient feedback analysis.
-Additionally, we compare their results with kBERT to assess its potential in addressing the challenges posed by limited and short healthcare-related textual data.
-These methods allow for a more comprehensive analysis of patient feedback, addressing both latent topic discovery and document similarity-based clustering to improve topic coherence and interpretability.
-Given the challenges associated with short-text length, small dataset sizes, and domain-specific issues in patient feedback, this research aims to explore the following research questions (RQs):
+This study presents a system for analyzing patient survey feedback from hospital visits, with an emphasis on identifying complaint-related topics.
+To address the challenges of short, sparse, and domain-specific text, we use a curated keyword list to filter negative feedback and apply unsupervised methods to uncover dominant themes.
+We evaluate traditional topic models (LDA, GSDMM) and embedding-based approaches (BERTopic, kBERT) on a small short-text dataset.
+By combining domain-aware filtering with neural embeddings, our approach aims to improve topic coherence and interpretability in low-resource healthcare settings.
+This work is guided by the following research questions (RQs):
 RQ1: How can unsupervised techniques be effectively utilized to identify and categorize topics in short-text patient feedback?
@@
 RQ3: How can the performance of unsupervised topic modeling and clustering methods be effectively evaluated in short-text patient feedback analysis?
-The rest of the paper is structured as follows: Section II reviews related work on complaint identification, topic modeling, and clustering, highlighting the challenges of short-text analysis in healthcare.
-Section III details the proposed methodology, including data preprocessing, keyword filtering, and applying unsupervised techniques, addressing RQ1.
-Section IV presents the result analysis, comparing the performance of both topic modeling and clustering approaches, evaluating their coherence, diversity, and interpretability, thereby addressing RQ3.
-Section V discusses the findings and their implications for healthcare service improvement, effectively answering RQ2.
-Finally, Section VI concludes the study.
 II Related Work
 II-A Topic Modeling and Word Embeddings-based Clustering
-In the field of NLP, complaint identification, particularly when labeled datasets are available, is often approached as a classification task.
-This area employs various machine learning models like support vector classifier (SVC), Random Forest, Decision Tree, and multilayer perception (MLP), along with deep learning classifiers such as Dense Neural Networks and convolutional neural networks (CNNs) [6].
-Additionally, NLP libraries like VADER, TextBlob, and NLTK are commonly used to derive sentiment scores, enabling the classification of feedback into positive, negative, or neutral categories [7].
-However, these libraries might not always capture nuanced details, particularly in domain-specific contexts.
-With the rise of Large Language Models (LLMs), the field of complaint classification is evolving.
-L.
-Yuan et al.
-introduced a Complaint Text Classification Framework (CCF) based on LLMs, which leverages advanced feature extraction techniques from large language models, significantly enhancing classification accuracy [8].
-While classification alone is useful, it doesn’t provide deep insights into the nature of the text.
-LDA can be effective in understanding the underlying concepts of language.
-LDA identifies the latent topics within a text corpus.
-As demonstrated in subsequent studies, LDA is frequently used for sentiment analysis in complaint data.
-For instance, A.
-Karami et al.
-utilized LDA to categorize GEICO customer complaints into four distinct groups [9], while N.
-Hu et al.
-applied LDA and Structural Topic Modeling to analyze hotel reviews in New York City, identifying ten major themes of customer dissatisfaction and how these vary across different hotel grades [10].
-M.
-N.
-Aziz et al.
-adopted a mixed approach for opinion classification, leveraging both Lexicon-Based (SentiWordNet) and Machine Learning-Based (TF-IDF with SVM) methods, and employed LDA for topic clustering [11].
-As research in this field progresses, various topic modeling methods have been compared to evaluate their effectiveness.
-For instance, H.
-S.
-Jung et al.
-evaluated four topic modeling methods—LDA, Nonnegative Matrix Factorization (NMF), Combined Topic Models (CTM), and BERTopic.
-Their findings showed that BERT based model demonstrated superior performance across both topic diversity and coherence [12].
-Embeddingbased methods have significantly advanced topic modeling by utilizing contextual embeddings like BERT and SBERT to capture semantic and syntactic nuances.
+In NLP, complaint identification is often treated as a classification task when labeled data is available.
+Traditional approaches use machine learning models such as SVC, Random Forest, and MLP, as well as deep learning methods like CNNs.
+While basic sentiment analysis tools (e.g., VADER, TextBlob, NLTK) can classify feedback as positive, negative, or neutral, they often miss domain-specific nuances.
+Recent advancements in large language models (LLMs) have significantly improved various NLP tasks [16].
+Frameworks like the Complaint Text Classification Framework (CCF) leverage LLM-based feature extraction to enhance classification accuracy [23].
+While classification is helpful, it often lacks deeper interpretability.
+Topic modeling techniques like LDA are widely used to uncover latent themes in textual data and have been applied across domains to analyze customer feedback.
+As the field evolves, several studies have compared topic modeling methods to assess their effectiveness.
+For example, Jung et al.
+evaluated LDA, NMF, CTM, and BERTopic, finding that BERT-based models outperformed others in both topic diversity and coherence [12].
+Embedding-based methods have significantly advanced topic modeling by utilizing contextual embeddings like BERT and SBERT to capture semantic and syntactic nuances.
 The Embedded Topic Model (ETM) [9] integrates LDA with embeddings to enhance topic coherence.
-Approaches combining BERT embeddings with clustering methods [14, 15] show improved coherence and interpretability.
-For instance, a framework integrating BERT with clustering and dimensionality reduction [14] demonstrated superior clustering quality, while a hybrid approach combining BERT, LDA [15] achieved high topic coherence.
-C-Top2Vec [16] further refined topic modeling through multi-vector representations and hierarchical clustering for granular insights.
-Topic modeling has evolved with the introduction of advanced models, yet traditional methods still hold significant advantages in certain applications.
-GPTopic, a novel approach leveraging LLMs, enhances the interpretability and interaction of topic representations, surpassing the traditional “top-word” approach by allowing dynamic modifications and chat-based interactions with topics.
-However, GPTopic can occasionally generate inaccurate results due to hallucinations, particularly when applied to datasets with fewer than 10,000 documents [17].
-While LLMs, including those used in GPTopic, offer flexibility and zero-shot capabilities, they struggle with maintaining topic granularity and often produce overlapping or redundant topics, which complicates the identification of distinct themes [18].
-In contrast, traditional methods like LDA and BERT base model remain consistent and reliable, generating coherent topics without hallucinations, making them highly adaptable across various domains without extensive fine-tuning [19].
-However, despite their advancements, LLMs require substantial computational resources, making their deployment costly due to the need for high-performance GPUs or TPUs for fine-tuning and inference.
-This makes traditional methods like LDA and clustering-based approaches more practical for resource-constrained environments.
-Additionally, LLM-enhanced applications that provide APIs or interfaces like ChatGPT [20], pose risks of data leakage, as models trained on large datasets can unintentionally expose sensitive information.
-The lack of clarity in how data is stored, processed, and transmitted further exacerbates trust, security, and privacy risks [21, 22].
-TABLE I: Datasets
-Dataset Name Size Median Type Availability
-Patient Feedback 439 14 Short-text private
-Clickbait-title 10,000 9 Short-text public
-Clickbait-title (500) 500 9 Short-text public
-20Newsgroup 15,021 83 Long-text public
-20Newsgroup (500) 500 86 Long-text public
 II-B Topic Modeling for Short-Text
-Conventional methods like LDA encounter challenges due to their reliance on bag-of-words representations, which can be inadequate for capturing the nuances in short texts[23].
-Recent research highlights the effectiveness of alternative models.
-For instance, GSDMM and BERTopic have shown promising results, outperforming LDA in handling short-text.
-One study by Udupa et al.
-[23] specifically explored the use of GSDMM and BERTopic for efficient topic clustering in short-text scenarios.
-Meanwhile, Jiang et al.
-[24] implemented the Associated Topic Model (ATM) which leverages a collapsed Gibbs sampling algorithm, further emphasizing the diversity of approaches suitable for short-text analysis.
-Additionally, clustering token-level contextualized word representations from models like BERT has proven effective in capturing polysemy and organizing documents, performing comparably or better than LDA topic models, as found in recent evaluations [25].
-These studies collectively suggest a shift towards models that can better interpret the limited contextual information available in shorter documents.
+Conventional methods like LDA struggle with short texts due to their reliance on bag-of-words representations, which often fail to capture contextual nuances [21].
+Recent studies have shown that alternative models like GSDMM and BERTopic outperform LDA in short-text scenarios.
+For example, Udupa et al.
+demonstrated the effectiveness of both GSDMM and BERTopic for topic clustering in brief texts [21].
+Additionally, clustering token-level contextualized embeddings from models like BERT has proven effective in capturing polysemy and organizing documents, often matching or exceeding LDA in performance [20].
+These findings indicate a shift toward models better suited for interpreting limited-context data.
 II-C Topic Modeling in Healthcare Feedback Service
@@
 To best of our knowledge, no prior study has addressed short-text topic modeling on limited datasets, specifically in the context of healthcare service.
-TABLE II: Sample Keyword (Hot Word) List
-infect angry malpractice delay MRSA
-blood sepsis cops grievance mis
-trauma wound abus argu un
-negl viol massac hipaa hippa
-african american gun bruise burn
+TABLE I: Sample Keyword (Hot Word) List
+Type Examples
+Full angry malpractice delay grievance african
+Short negl abus argu negl Viol
+Jargon hipaa hippa MRSA
+Prefix un mis
 III Methodology
-This study focuses on short-text patient feedback, specifically in small datasets, where traditional topic modeling techniques often face challenges.
-We apply keyword filtering to identify complaint-related feedback using a curated lexicon, ensuring a focused analysis on patient grievances.
-Simultaneously, topic modeling is employed to extract dominant themes, providing insights into both complaints and general concerns.
-As our dataset consists of short-text feedback, GSDMM is used as the baseline for short-text topic modeling, while LDA is included as a reference for long-text topic modeling.
-These models are compared with BERTopic and kBERT, which leverage transformer-based embeddings to enhance topic coherence in small short-text datasets.
+This study analyzes short-text patient feedback using keyword filtering and unsupervised topic modeling.
+GSDMM serves as the baseline for short-text modeling, while LDA provides a long-text reference.
+We also implement BERTopic and kBERT, which leverage transformer-based embeddings to enhance coherence.
 III-A Dataset Description
@@
 The dataset is private and consists of short-text feedback, with a median word count of 14.
-To evaluate the robustness of our methodology, we also tested it on two additional public datasets: the Clickbait-title dataset [29] and the 20Newsgroup dataset[30].
-The Clickbait-title dataset is a short-text dataset comprising headlines categorized as clickbait or non-clickbait.
-We selected 10,000 samples from this dataset, with a median word count of 9, making it representative of short-text content.
-Additionally, to examine model performance on small short-text datasets, we extracted a subset of 500 samples (clickbait-title (500)).
-In contrast, the 20Newsgroup dataset consists of longer articles from 20 distinct topics.
-It is a widely used benchmark dataset for different natural language tasks.
-To standardize the data, we excluded the bottom and top 10% of articles based on word count and retained 15,021 articles for evaluation and the median word count is 83.
-We also extracted a subset of 500 samples (20Newsgroup (500)) to explore small dataset scenarios in long-text settings.
+To evaluate the robustness of our methodology, we also tested it on two public datasets: the Clickbait-title dataset [6] and the 20Newsgroup dataset [1].
+The Clickbait-title dataset contains short headlines labeled as clickbait or non-clickbait.
+We selected 10,000 samples with a median word count of 9, and also created a smaller subset of 500 samples to evaluate performance on small short-text data.
+The 20Newsgroup dataset consists of longer articles across 20 distinct topics and is widely used for benchmarking NLP tasks.
+A 500-sample subset was extracted to simulate small long-text scenarios.
 These datasets allow for a comprehensive evaluation of topic modeling and clustering techniques across varying text lengths (short vs.
@@
 III-B Keyword Processing
-In this work, we utilize a carefully curated keyword list to effectively filter and categorize the patient feedback into positive and negative comments.
-These keywords, which are crucial for analyzing the sentiment of the feedback, are presented in Table II.
-This lexicon ensures a comprehensive reflection of the terminology commonly found in patient complaints, aiding in the accurate assessment of their sentiment.
-The keyword list is also important as it contains domain-specific nuances.
-Some feedback includes medical and legal terms that may not explicitly mention complaints.
-For instance, “I saw another patient’s information on the clinic’s screen, which seems like a HIPAA issue”.
-These subtle details are often missed by general-purpose models.
-Capturing these nuances is crucial for improving patient care.
-In Table II, the keywords are systematically classified into four distinct groups to enhance the precision of complaint detection in the texts.
-This classification is as follows:
-  • Full Keywords: These encompass complete terms that are often used in their entirety within complaints, such as “accident”, “delay”, “agitate”, “american” etc.
-  • Shortened Keywords: Truncated forms of words, like “abus” (for abuse) and “argu” (for argue), which are used in different parts of speech (POS) in the feedback.
-  • Prefix Keywords: Terms like “mis” and “un”, serve as prefixes, indicative of negative connotations when attached to other words.
-  • Jargon/Medical Term Keywords: This category includes essential terms like “sepsis”, “bruise”, “HIPAA”, and “MRSA”, crucial for identifying comments on various topics such as regulatory compliance or infection-related concerns.
-‘HIPPA’ is included as it is a frequent misspelling of ‘HIPAA’ by patients.
-Each group is designed to address the different ways keywords could manifest in the data — as standalone words, abbreviated forms, or as part of medical jargon — thus providing a comprehensive filter for the preliminary categorization of comments.
-In our methodology, the keyword processing phase presented a unique challenge — the ambiguity arising from shortened keywords.
-For example, the abbreviation ‘compl’ is extracted from terms such as ‘complaint’ and ‘complete’, though only the former is relevant to our analysis.
-To mitigate this, we refined our keyword list to include only the full, unambiguous forms of each word, ensuring that only contextually relevant terms are considered.
-Another complexity arose from the morphological forms of keywords within the dataset.
-Words like ‘argue’ appeared in various derivatives, such as ‘argues’, ‘argued’, or ‘arguing’.
-Traditional lemmatization is employed to address this, which involves reducing words to their base or dictionary form.
-Lemmatization’s effectiveness is contingent upon accurate POS tagging.
-The process needs to recognize the grammatical role of words within the context of their use.
-This is because POS tagger models are typically trained on complete sentences or documents rather than isolated tokens.
-This training approach poses a challenge when a word like ‘argument’ needs to be reduced to ‘argue’ [31].
-To navigate these linguistic nuances, we utilize the ‘word_forms’ library in Python [32].
-This library provides a comprehensive mapping of word forms across different POS categories: nouns, verbs, adverbs, and adjectives; allowing us to accurately identify and process the various permutations of a keyword in our dataset.
+In this work, we utilize a carefully curated list of domain-specific keywords to filter patient feedback into complaint and non-complaint categories.
+These keywords capture subtle indicators of dissatisfaction, such as legal, medical, or emotional terms that are often overlooked by general-purpose models.
+We grouped the keywords into four categories: full, shortened, prefix, and jargon.
+These categories, along with representative examples, are shown in Table I.
+For instance, a comment like “I saw another patient’s information on the clinic’s screen, which seems like a HIPAA issue” highlights a potential privacy concern without explicitly using negative sentiment words, which domain-specific filtering can effectively capture.
+A key challenge in keyword processing is handling ambiguous terms, especially shortened forms.
+For example, the abbreviation compl could represent either complaint or complete, but only the former is relevant in our context.
+To avoid such ambiguity, we refined the list to include only unambiguous and contextually meaningful terms.
+Another complexity arises from morphological variations of keywords.
+Words like argue may appear as argues, argued, or arguing in the dataset.
+Addressing this requires effective lemmatization, which reduces words to their base forms.
+However, the success of lemmatization depends heavily on accurate part-of-speech (POS) tagging.
+Most POS taggers are trained on full sentences rather than isolated tokens, making context-sensitive lemmatization particularly challenging [3].
+To manage these linguistic nuances, we employ the word_forms Python library.
+This tool provides comprehensive mappings of words across POS categories—nouns, verbs, adverbs, and adjectives—allowing us to robustly identify and process all morphological variants of keywords in our dataset.
 The complete keyword processing procedure is detailed in Algorithm 1.
@@
 III-C Feedback Preprocessing
-The fidelity of topic modeling in NLP is heavily reliant on the quality of input data.
-Recognizing this, our methodology is built on a thorough protocol for preprocessing data.
-Initially, tokenization—the segmentation of text into individual terms—is conducted to granulate the dataset into analyzable units.
-Subsequent to tokenization, we focus on the elimination of stop words.
-These words, while structurally integral, contribute minimally to the semantic load of the text.
-Common stop words such as ‘a’, ‘and’, ‘the’, and ‘of’ are identified and removed.
-Their high frequency and low informational value make them prime candidates for exclusion to enhance data purity.
-To perform this task, we employ the ‘spacy’ library—a powerful and versatile Python toolkit for advanced NLP [33].
-Its sophisticated algorithms allowed for an efficient and thorough filtration process, excising stop words from the tokenized text.
-This cleansing step is vital, as it significantly reduces noise within the data, thereby streamlining the path for more accurate and meaningful topic modeling.
-The detailed procedure for this filtration is presented in Algorithm 2.
+The effectiveness of topic modeling in NLP depends heavily on the quality of input data.
+To ensure this, our methodology follows a structured preprocessing pipeline.
+We begin with tokenization, which segments text into individual terms to make it analyzable.
+Next, we remove stop words such as “a,” “and,” “the,” and “of,” which are frequent but carry little semantic value.
+This step helps reduce noise and improve data clarity.
+We use the “spacy” library for stop word removal due to its robust and efficient NLP capabilities.
+This preprocessing step is essential for producing cleaner input that enhances the accuracy and interpretability of downstream topic modeling.
+The detailed steps are outlined in Algorithm 2.
 III-D Topic Identification
 III-D1 Latent Dirichlet Allocation (LDA)
-LDA is a generative probabilistic model widely used for uncovering latent topics within a collection of text documents [34].
-LDA assumes that each document is a mixture of various topics, where each topic is characterized by a probability distribution over words.
-By breaking down a document-word matrix into two smaller matrices—the Document-Topic Matrix and the Topic-Word Matrix—LDA reveals the relationships between documents and their underlying topics, as well as the association between topics and the defining words.
-This process makes LDA particularly valuable for analyzing large collections of text data.
-The LDA model operates through a hierarchical generative process.
-At the document level, a topic distribution $\theta$ for each document $d$ is drawn from a Dirichlet distribution with parameter $\alpha$, which serves as the prior for the document-topic distributions.
-For each word $w_{d,n}$ in a document, a topic $z_{d,n}$ is sampled from the document’s topic distribution $\theta_{d}$.
-Subsequently, a word is generated by sampling from the topic-specific word distribution $\beta_{z_{d,n}}$, which is drawn from another Dirichlet distribution with parameter $\eta$.
-This layered structure captures both the document-level topic diversity and the word-level thematic coherence.
-Figure 1: LDA Blueprint (Source:[13, 35] )
-In our study, we applied LDA to analyze hospital patient feedback documents, identifying and categorizing latent themes within the text.
-Each document was modeled as a probabilistic mixture of topics, with $\theta_{d,k}$ representing the likelihood of topic $k$ in document $d$, and $\beta_{k,w}$ indicating the relevance of word $w$ within topic $k$.
-To enhance thematic clarity, we experimented with 3 to 20 topic clusters to balance specificity and breadth.
-The Dirichlet parameter $\alpha$ was set to ‘auto,’ allowing dynamic adjustment of document-topic distributions.
-We also fine-tuned the top words per topic, improving the model’s ability to highlight relevant terms and themes.
+LDA is a generative probabilistic model that uncovers latent topics within a document collection by modeling each document as a mixture of topics, where each topic is a distribution over words [4].
+It decomposes the document-word matrix into a document-topic and a topic-word matrix, revealing underlying thematic structures.
+At the core of LDA, topic proportions for each document are drawn from a Dirichlet prior, and words are generated based on topic assignments and word distributions.
+In our study, we applied LDA to hospital patient feedback to identify latent themes.
+We experimented with 3 to 20 topic clusters to balance specificity and generality, using the auto setting for Dirichlet priors to adapt topic distributions.
+Top words per topic were fine-tuned to enhance interpretability.
 III-D2 Gibbs Sampling Dirichlet Mixture Model (GSDMM)
@@
 The implementation of GSDMM involved modeling each document as a mixture of topics, with a single dominant topic assigned to each document.
-For a given document $d$, the topic $k$ is assigned based on the conditional probability $P(k\mid d)$, which is calculated using Equation 1.
+For a given document $d$, the topic $k$ is assigned based on the conditional probability P(kd)P(k∣d), which is calculated using Equation 1.
 In this equation, $n_{k,-d}$ represents the number of documents assigned to topic $k$, excluding the current document $d$; $N_{-d}$ is the total number of documents excluding $d$; $K$ is the total number of topics; is the count of word $n_{k,w}$ w in topic k; $n_{k}$ is the total number of words assigned to topic $k$; and $V$ is the vocabulary size.
@@
 III-D3 BERTopic
-BERTopic [38] is a topic modeling technique that integrates transformer-based embeddings, dimensionality reduction, clustering, and topic representation to extract meaningful insights from textual data.
-It begins with a transformer embedding model, such as BERT or RoBERTa, to generate high-dimensional contextual embeddings that capture semantic relationships between words and documents.
-To enhance computational efficiency and clustering performance, UMAP (Uniform Manifold Approximation and Projection) is applied for dimensionality reduction, preserving essential structures in a lower-dimensional space.
-The reduced embeddings are then clustered using HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise), which dynamically determines the optimal number of topics based on data density, enabling the discovery of latent topics in an unsupervised manner.
-This method allows BERTopic to automatically generate topics without requiring the user to specify a predefined number.
-Additionally, BERTopic framework [39] offers the flexibility to manually specify the desired number of topics, allowing for both unsupervised and user-controlled topic generation.
-Finally, c-TF-IDF (class-based Term Frequency-Inverse Document Frequency) is employed to extract the most representative words for each cluster, improving topic interpretability.
-Combining these four components, BERTopic provides a scalable and flexible approach to topic modeling, making it particularly effective for analyzing large, unstructured text data.
-However, BERTopic has some limitations.
-It assumes each document belongs to a single topic, which may not fully capture multi-topic documents.
+BERTopic [10] is a topic modeling technique that combines transformer-based embeddings, dimensionality reduction, clustering, and topic representation to extract insights from text.
+It uses models like BERT or RoBERTa to generate contextual embeddings, which are reduced in dimensionality using UMAP for efficient clustering.
+HDBSCAN is then applied to group similar texts and determine the number of topics based on data density.
+BERTopic also allows users to set a fixed number of topics manually.
+For interpretability, it employs class-based TF-IDF (c-TF-IDF) to extract representative words per cluster.
+While scalable and flexible, BERTopic assumes each document belongs to a single topic, which may limit its effectiveness on multi-topic texts.
 III-D4 k-means clustering with BERT embedding (kBERT)
-To address the unique challenges posed by our small dataset of short-texts, we developed a customized topic modeling pipeline that integrates a customized pre-processing approach, BERT embeddings and k-means clustering.
-Our topic identification system leverages the Bidirectional Encoder Representations from Transformers (BERT), a transformer-based model pre-trained on a vast corpus of text, to generate rich, contextualized embeddings.
-These embeddings are particularly effective in capturing the subtle nuances of short-text data, making them ideal for our application.
-Additionally, BERT embeddings can effectively handle Out-of-Vocabulary (OOV) words, making them particularly useful for dealing with domain-specific terminology.
-Importantly, we only utilized the embedding layer of BERT while freezing all other layers.
+To address the unique challenges posed by our small dataset of short-texts, we developed a customized topic modeling pipeline that integrates a customized pre-processing approach, BERT embeddings, and k-means clustering.
+BERT, a transformer-based model pre-trained on a large corpus, generates rich contextual embeddings that capture subtle linguistic nuances and effectively handle Out-of-Vocabulary (OOV) terms, particularly useful for domain-specific language.
+We only utilized the embedding layer of BERT while freezing all other layers.
 This ensured that the model acted purely as a feature extractor, preventing any modifications to its pre-trained weights.
@@
 The high-dimensional embeddings generated by BERT are preserved to retain the full contextual representation.
-Our approach began with careful preprocessing of the textual data.
-Stopwords were removed, and the text was lemmatized to ensure uniformity.
-Additionally, we extended the stopword list with domain-specific and custom terms to improve data quality.
-Each preprocessed comment was then encoded into embeddings using the ‘bert-base-uncased’ [^1] https://huggingface.co/google-bert/bert-base-uncased variant of the pre-trained BERT model.
-This model was chosen for its balance between performance and computational efficiency, as well as its ability to generalize well across diverse datasets.
-The uncased variant ensured the focus remained on content, minimizing case-sensitive variations, which is particularly advantageous for short-texts.
-The generated embeddings served as input for a k-means clustering algorithm, where we experimented with different numbers of clusters ($n$ = 3, 5, 10 etc) to find the optimal grouping for the data.
-For topic interpretability, we extracted the top $k$ words (e.g., 5,10) from each cluster by analyzing word frequencies in representative comments.
-To improve topic interpretability, we first selected a set of representative comments for each cluster based on their proximity to the cluster centroid in the BERT embedding space.
-These comments were then analyzed to extract the most frequently occurring words, ensuring that the topic words accurately reflected the cluster’s semantic meaning.
-This iterative process allowed us to refine the topic selection and ensure the clusters were meaningful and cohesive.
+Our approach began with careful preprocessing of the textual data, including stopword removal and lemmatization for consistency.
+We also extended the stopword list with domain-specific terms to improve data quality.
+Each comment was then encoded using the ‘bert-base-uncased’ [^1] https://huggingface.co/google-bert/bert-base-uncased model, selected for its balance of performance and generalization.
+Using the uncased variant helped reduce case-related variability, which is beneficial for short texts.
+The resulting embeddings were clustered using k-means, with various cluster sizes ($n$ = 3, 5, 10, etc.) tested to find optimal groupings.
+For interpretability, we identified representative comments closest to each cluster centroid and extracted the most frequent words from them.
+These words formed the topic labels, ensuring alignment with the semantic meaning of each cluster.
+This iterative process refined topic selection and enhanced cluster coherence.
 IV Result Analysis
@@
 This increment underscores the pivotal role of nuanced linguistic techniques in enhancing the sensitivity of our keyword-based feedback-filtering approach.
-[…]
+TABLE II: Coherence Score ($C\textsubscript{v}$) on Large Datasets
 Dataset Model Number of Topics
@@
 kBERT 0.36 0.39 0.40 0.47 0.43
-IV-B Topic Modeling
-In their work, R.
-Egger et al.
-highlight the importance of selecting the appropriate model based on the nature of the data and emphasize the need for careful interpretation of results [42].
-Building on this perspective, our topic modeling efforts focused on finding the ideal balance between coherence and clarity.
-To ensure a comprehensive evaluation, we assessed our models’ performance using two metrics: the widely used coherence score ($C_{v}$) for topic interpretability and Rank-Biased Overlap (RBO) for evaluating topic diversity.
-Coherence score $C_{v}$ quantitatively assesses the coherence of topics.
-A higher $C_{v}$ score indicates better performance, suggesting that the top words of a topic are not only frequent but also semantically similar.
-This indicates a clearer and more meaningful topic representation.
-The $C_{v}$ measure is derived from the Normalized Pointwise Mutual Information (NPMI) [43].
-It accounts for the probability of words co-occurring within topics relative to their individual probabilities.
-The NPMI tends to be particularly attuned to less common words and is well-suited for analyzing small datasets like ours [44].
-It can be calculated using the equation 2 [45].
-In this equation, $\mathrm{P}\left(w_{i}^{r},w_{i}^{s}\right)$ is the joint probability of the words occurring together, while $\mathrm{P}\left(w_{i}^{r}\right)$ and $\mathrm{P}\left(w_{i}^{s}\right)$ are the individual probabilities of the words.
-The epsilon $\epsilon$ is a small number added to prevent division by zero.
-The $C_{v}$ is defined by equation 3.
-$\gamma$ is a parameter that can be tuned to give more importance to higher PMI values.
-This tuning allows for a weighted contribution of word pairs based on their PMI.
-$\operatorname{NPMI}\left(w_{i}^{r},w_{i}^{s}\right)=\frac{\log_{2}\frac{P\left(w_{i}^{r},w_{i}^{s}\right)+\epsilon}{P\left(w_{i}^{r}\right)P\left(w_{i}^{s}\right)}}{-\log_{2}\left(P\left(w_{i}^{r},w_{i}^{s}\right)+\epsilon\right)}$ (2)
-$\mathrm{C}_{\mathrm{V}}\left(w_{i}^{r},w_{i}^{s}\right)=\mathrm{NPMI}^{\gamma}\left(w_{i}^{r},w_{i}^{s}\right)$ (3)
-To evaluate the effectiveness of different topic modeling approaches, we computed the Coherence Score (Cv) for each model across three datasets: Clickbait-title, 20Newsgroup, and Patient Feedback.
-The datasets were divided into large and small subsets to assess model robustness across different data scales.
-Given that the Patient Feedback dataset consists of short-text comments, we tested our models on public short-text datasets for validation.
-For the analysis, we primarily focus on short-text, especially for small datasets, to evaluate how well our models handle limited data.
-Coherence scores are measured using the top 10 topic words for large-text datasets and the top 5 topic words for short-text datasets.
-The best results for short-text are marked boldface in Tables III and IV.
-For large datasets, we generated topics at different levels (5, 10, 20, 30, and 40 topics).
-As shown in Table III, GSDMM consistently outperformed other models in short-text datasets, achieving the highest coherence score of 0.49 for 20 topics in Clickbait-title.
-This indicates its ability to effectively capture coherent topics in short-texts.
-Additionally, kBERT also achieved a high coherence score of 0.49 at 30 topics for short-text datasets (Clickbait-title).
-This motivates further exploration of kBERT for small datasets, where capturing meaningful patterns in limited data is crucial.
-LDA, while performing well in long-text scenarios, struggled in short-text datasets.
-BERTopic, which dynamically determines the number of topics, generated 163 topics in 20Newsgroup (Cv = 0.60) and 178 topics in Clickbait-title (Cv = 0.49).
-This adaptability allowed BERTopic to excel in large datasets by identifying a diverse range of topics.
-However, in smaller datasets, it generated a limited number of topics, reducing coherence performance.
-In the Patient Feedback dataset, BERTopic produced 3 topics with a coherence score of 0.46, indicating moderate effectiveness in short-text scenarios.
-For small datasets, we generated 3, 5, 10, 15, and 20 topics.
-As presented in Table IV, kBERT performed exceptionally well in handling short-text patient feedback, achieving the highest coherence score of 0.53 for 3 topics, demonstrating its strong ability to extract meaningful insights from limited and sparse patient feedback data.
-Additionally, in the Clickbait-title (500) dataset, kBERT achieved a strong coherence score of 0.45 for 5 topics, reinforcing its strength in short-text small dataset scenarios.
-This suggests that BERT embeddings are particularly useful for extracting meaningful insights from short and sparse healthcare-related feedback.
-[…]
+a generated 178 topics ($C_{v}$=0.48), b generated 163 topics ($C_{v}$=0.60)
+TABLE III: Coherence Score ($C\textsubscript{v}$) on Small Datasets
 Dataset Model Number of Topics
@@
 GSDMM 0.49 0.43 0.42 0.46 0.46
-BERTopicc 0.46 N/A N/A N/A N/A
+BERTopic 0.46 N/A N/A N/A N/A
 kBERT 0.53 0.49 0.48 0.48 0.41
-TABLE V: Topics from kBERT with Examples
-Topic Name Topic words Example Complain?
-Staff Conduct and Cleanliness good, clean, great, everyone, professional Everyone was careful in the emergency department No
-Safety Protocols compliance mask, staff, wearing, safe, one The staff’s masks were reassuring in these uncertain times. No
-Efficiency of Service doctor, time, appointment, waiting, room Very long wait from the time I entered the room until the doctor arrived Yes
-TABLE VI: IRBOavg Score on Top 5 words and 3 Topics
+a generated 2 topics ($C_{v}$=0.40), b generated 2 topics ($C_{v}$=0.44)
+IV-B Topic Modeling
+We aimed to balance coherence and clarity in topic modeling.
+To evaluate performance, we used two metrics: Coherence score ($C_{v}$) for interpretability and Rank-Biased Overlap (RBO) for topic diversity.
+Coherence score ($C_{v}$) quantitatively measures how semantically related the top words in a topic are.
+A higher $C_{v}$ indicates better topic quality, suggesting that the most representative words frequently co-occur and are contextually meaningful.
+The $C_{v}$ score is derived from Normalized Pointwise Mutual Information (NPMI) [5], which considers the joint and individual probabilities of word co-occurrence.
+NPMI is defined in Equation 2, where $\epsilon$ prevents division by zero.
+The final $C_{v}$ score is calculated as shown in Equation 3, where $\gamma$ adjusts the weight given to high-PMI word pairs.
+$\operatorname{NPMI}\left(w_{i}^{r},w_{i}^{s}\right)=\frac{\log_{2}\frac{P\left(w_{i}^{r},w_{i}^{s}\right)+\epsilon}{P\left(w_{i}^{r}\right)P\left(w_{i}^{s}\right)}}{-\log_{2}\left(P\left(w_{i}^{r},w_{i}^{s}\right)+\epsilon\right)}$ (2)
+$\mathrm{C}_{\mathrm{V}}\left(w_{i}^{r},w_{i}^{s}\right)=\mathrm{NPMI}^{\gamma}\left(w_{i}^{r},w_{i}^{s}\right)$ (3)
+To assess model effectiveness, we computed the $C_{v}$ coherence score for each method across three datasets: Clickbait-title, 20Newsgroup, and Patient Feedback.
+These were split into large and small subsets to test robustness at different data scales.
+As Patient Feedback consists of short-text comments, we included public short-text datasets for validation.
+Coherence was measured using the top 10 topic words for long texts and the top 5 for short texts.
+Best-performing short-text results are highlighted in Tables II and III.
+For large datasets, we evaluated models across multiple topic counts (5 to 40).
+As shown in Table II, GSDMM consistently achieved high coherence on short-text data, scoring 0.49 for 20 topics in Clickbait-title.
+kBERT also performed well, reaching a score of 0.49 at 30 topics, highlighting its strength in small, short-text scenarios.
+In contrast, LDA struggled with short texts but performed better on long documents.
+BERTopic, which automatically determines the number of topics, generated 163 topics in 20Newsgroup ($C_{v}$ = 0.60) and 178 in Clickbait-title ($C_{v}$ = 0.49), showing strong performance on large datasets.
+However, its coherence dropped in smaller datasets, producing only 3 topics ($C_{v}$ = 0.46) in Patient Feedback.
+For small datasets, we generated 3, 5, 10, 15, and 20 topics.
+As shown in Table III, kBERT achieved the highest coherence score of 0.53 for 3 topics on Patient Feedback, demonstrating its effectiveness in short and sparse data.
+It also performed well on the Clickbait-title (500) dataset, scoring 0.45 for 5 topics, reinforcing the usefulness of BERT embeddings in low-resource short-text scenarios.
+Rank-Biased Overlap (RBO) is a similarity measure for comparing ranked lists, placing more emphasis on top-ranked items and handling incomplete or non-conjoint lists [22].
+It is particularly useful for evaluating topic diversity by assessing overlap among top topic words.
+RBO ranges from 0 (no overlap) to 1 (identical lists), with the parameter $p$ controlling the depth of comparison.
+To better capture topic diversity, we use the inverted RBO $(1-\operatorname{RBO})$, where higher values indicate more distinct topics.
+We compute the average inverted RBO across all topic pairs, as shown in Equations 4 and 5.
+$\operatorname{RBO}(S,T,p)=(1-p)\sum_{d=1}^{\infty}p^{d-1}\cdot A_{d}$ (4)
+$\begin{split}\operatorname{IRBO_{avg}}(A,B,C,p)=\frac{1}{3}\Big[(1-\operatorname{RBO}(A,B,p))+\\ (1-\operatorname{RBO}(A,C,p))+\\ (1-\operatorname{RBO}(B,C,p))\Big]\end{split}$ (5)
+TABLE IV: IRBOavg Score on Top 5 words and 3 Topics
 Dataset Model Score
@@
 kBERT 1.00
-Rank-Biased Overlap (RBO) is a similarity measure for ranked lists that emphasizes top-weightedness and accommodates incomplete or non-conjoint lists [46].
-It is particularly useful for evaluating topic diversity in topic modeling by comparing the top-ranked terms across topics.
-The RBO is defined by the equation 4.
-$A_{d}$ is the agreement at rank $d$, and $p$ ($10<p<$ 1) adjusts the weight of deeper ranks.
-Higher $p$ values increase the depth considered, ensuring meaningful topic comparisons.
-RBO scores range from 0 (no overlap) to 1 (identical lists).
-Adjust the parameter $p$ to focus on higher or deeper ranks.
-To provide a more comprehensive diversity measure, we used Inverted RBO $(1-\operatorname{RBO})$, where higher values indicate greater topic diversity by reflecting lower overlap in top-ranked terms.
-We further compute the Averaged Inverted RBO by averaging pairwise inverted RBO values, as defined by equation 5.
-Here, A, B, and C represent the ranked lists of top words from three different topics, and their pairwise inverted RBO scores quantify the extent of dissimilarity among them.
-A higher $\operatorname{IRBO_{avg}}$ value indicates greater diversity in the extracted topics.
-$\operatorname{RBO}(S,T,p)=(1-p)\sum_{d=1}^{\infty}p^{d-1}\cdot A_{d}$ (4)
-$\begin{split}\operatorname{IRBO_{avg}}(A,B,C,p)=\frac{1}{3}\Big{[}(1-\operatorname{RBO}(A,B,p))+\\ (1-\operatorname{RBO}(A,C,p))+\\ (1-\operatorname{RBO}(B,C,p))\Big{]}\end{split}$ (5)
-In this study, we selected 3 topics for the Patient Feedback dataset and evaluated topic diversity using Average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$).
-A higher IRBO score indicates greater topic distinctiveness by measuring lower overlap in top-ranked terms.
-As shown in Table VI, BERTopic and kBERT achieved the highest diversity score (1.00), suggesting completely distinct topic distributions.
-GSDMM (0.97) also demonstrated strong topic separation, while LDA (0.83) exhibited comparatively lower diversity, indicating some overlap in identified topics.
+We evaluated topic diversity on the Patient Feedback dataset using Average Inverted Rank-Biased Overlap ($\operatorname{IRBO_{avg}}$), where higher scores indicate more distinct topics.
+As shown in Table IV, kBERT and BERTopic achieved perfect diversity scores of 1.00, while GSDMM scored 0.97 and LDA 0.83, indicating some topic overlap in LDA’s results.
 V Discussion
@@
 By applying a targeted keyword-based filtration method, we identified 102 complaint-related sentences out of 439 total comments.
-In evaluating the topic modeling approaches, kBERT consistently demonstrated superior performance in handling short-text complaint identification, achieving the highest coherence score ($C_{v}$ = 0.53 for 3 topics) and an IRBO score of 1.00, indicating distinct and well-separated topics.
-This suggests that BERT-based embeddings effectively capture the nuanced meaning of patient complaints, making it the most reliable model for healthcare-related short-text analysis.
-GSDMM also performed well, particularly in short-text datasets, as its cluster-based approach assigns a dominant topic to each document, making it effective for structured short-form feedback.
-BERTopic, while excelling in larger datasets due to its dynamic topic generation, struggled in smaller datasets, as observed in its moderate coherence score ($C_{v}$ = 0.46) for the Patient Feedback dataset.
-LDA, a traditional topic modeling approach, showed strong performance in long-text settings but struggled with short-text patient complaints, reaffirming its dependence on a larger corpus for meaningful topic extraction.
-These findings emphasize the importance of choosing context-aware models like kBERT for short-text analysis in healthcare settings.
-The combination of keyword filtering and effective topic modeling ensures that patient concerns are identified and addressed accurately, leading to data-driven improvements in service delivery.
-The insights gained from this study can help healthcare providers prioritize patient concerns, streamline complaint resolution, and enhance overall service efficiency, ultimately improving the patient experience and quality of care.
+TABLE V: Topics from kBERT with Examples
+Topic words Example Complain?
+good, clean, great, everyone, professional Everyone was careful in the emergency department No
+mask, staff, wearing, safe, one The staff’s masks were reassuring in these uncertain times. No
+doctor, time, appointment, waiting, room Long wait from the time I entered the room until the doctor arrived Yes
+In evaluating the topic modeling approaches, kBERT consistently outperformed other models in short-text complaint analysis, achieving the highest coherence score ($C_{v}$ = 0.53) and perfect topic diversity (IRBO = 1.00), indicating clear and well-separated topics.
+This highlights the strength of BERT embeddings in capturing the nuances of patient feedback.
+GSDMM also performed well, benefiting from its ability to assign a dominant topic per document.
+In contrast, BERTopic, while effective on large datasets, showed limited performance on smaller ones ($C_{v}$ = 0.46).
+LDA performed better on long texts but struggled with short, sparse data.
+These results underscore the value of context-aware models like kBERT in healthcare settings.
+Combined with keyword filtering, this approach enables accurate identification of patient concerns, supporting data-driven service improvements and enhanced care delivery.
 VI Conclusion
@@
 Additionally, a keyword-based filtering approach significantly improved complaint detection.
-The identified complaint categories—staff conduct and cleanliness, safety protocol compliance, and efficiency of service—provide actionable insights for healthcare service improvement.
-In the future, we will also explore few-shot and instruction-based prompting to improve complaint classification and topic extraction in healthcare feedback analysis.
-To achieve this, we will use open-source LLMs (e.g., LLaMA) locally with sufficient computational resources, preventing data leakage and ensuring compliance.
+The identified complaint categories include staff conduct and cleanliness, safety protocol compliance, and efficiency of service, and provide actionable insights for healthcare service improvement.
+Acknowledgment
+We gratefully acknowledge the support of the Advancing a Healthier Wisconsin (AHW) for funding this project.
+The authors used AI tools in this manuscript, including Grammarly for grammar checking and ChatGPT for rephrasing and refining, with careful review to ensure contextual accuracy.